Date Time MST Site Description Status
         
06/07/06 14:00 ARO I have posted the summer shutdown schedules for both telescopes here. News
06/07/06 11:00 SMT We have another two fires on Mt. Graham. Here is the latest news on these fires. News
06/06/06 12:30 12M I have posted procedures for shutting down the 12 Meter MAC system. They can be found here. News
06/06/06 10:30 ARO We continue to receive reports of a few observers who are having problems using Xremote. We believe the biggest problem is that the observer is starting their ssh session into the telescope computers without turning on X11 tunneling. Be sure to add the '-X' switch to your invocation of ssh. Complete documentation on Remote Observing can be found here. Hints
06/06/06 10:15 12M The 23mmsis receiver is being warmed up in an attempt to locate the problems with its cooling and maintaining a 4K temperature.  
06/05/06 16:00 12M MAC shutdown procedures are posted here. News
06/03/06 09:35 12M The 23mmsis just will not stay cold during the hottest parts of the day. An ongoing investigate continues.  
06/01/06 14:00 12M A small AC unit is installed in an attempt to cool the 23mmsis dewar and allow it to maintain its 4K JT temperature. News
05/30/06 12:00 12M The MAC is back online after a serious rebooting. Repaired
05/27/06 19:00 12M The MAC will not boot up properly and is shutdown until it can be repaired.  
05/24/06 14:00 ARO Director Lucy Ziurys is interviewed for the SETI Institute's Weekly Radio Program. Here is a shortened version containing only Lucy's segment on comets. News
05/22/06 06:45 MST Here is an update on the Mt. Graham "Taylor File". News
05/20/06 10:01 SMT The first fire of the 2006 fire season is burning on Mt. Graham. Here is a link to our 2006 fire season web page. News
05/18/06 14:30 12M The hot temperature are effecting more then just the receiver. Now the elevation encoder has a bit that appears to stick on whenever the temperature get above 34C. Investigation continues.  
04/28/06 11:45 ARO ARO now has a faq web page. Not much there yet, but it's a start. News
04/19/06 14:00 SMT The Forest Service has put out a news release on fire restrictions effective today, April 19, 2006 Please read the new release here. News
04/12/06 14:00 12M Stefanie Milam has submitted a CBET circular about her 12 Meter observations on the comet 73P. Here is the release:

Comet 73P/Schwassmann-Wachmann 3

S.N. Milam, A.J. Apponi, L.M. Ziurys, and S. Wyckoff, Arizona Radio Observatory and Steward Observatory

Observations of the J=1-0 transition of HCN at 88.6 GHz were conducted towards comet 73P component 'C' on April 11.36 UT with the Arizona Radio Observatory 12m telescope on Kitt Peak, AZ. Two out of three hyperfine components were detected with intensities of 0.035 +/- 0.005 K for the F=2-1 transition and 0.024 +/- 0.006 K for the F=1-1 line, with corresponding linewidths of 1.0 +/- 0.2 km/s and 0.7 +/- 0.2 km/s, respectively. These measurements indicate a total column density of ~3.4x10^{11} cm^{-2} for HCN in the 'C' source of 73P. A production rate of Q(HCN) ~ 3x10^{25} s^{-1} was calculated using a Monte Carlo model, assuming that this molecule is a parent species arising from the nucleus. This transition of HCN appears at the cometocentric velocity derived from the JPL Horizons ephemeris.

   Here is a link to the Harvard web site

News
02/14/06 20:00 12M 23mm receiver start cooling. Repaired
02/14/05 16:00 12M Turns out the brake was stuck engaged and not letting the dome door motor move. Once free of the brake, it moved freely. George ground a little clearance in the three tangs that engage in the frame of the motor. Once reassembled the motor worked fine. Replaced drive chains for both motors and tested. Dome door functions normally now. Repaired
02/13/06 16:00 12M Because the dome door problem look like it will take a few days to repair, we are warming up the receiver to try and unblock the JT stage.  
02/13/06 14:00 12M While switching from site generator back to commercial power, the 23mm receiver had a difficult time coming back down to operating temperature.  
02/13/06 8:00 12M The dome door is stuck in the open position and can not be closed. We are working on it.  
01/31/06 15:00 SMT We have now posted the design and repair manuals for the Forbes Filter Banks. (ARO Only) News
01/05/06 13:00 ARO We now have the entire weather history from both telescopes on-line and available here. News
01/04/06 14:00 ARO We have improved the 12 Meter weather page. See here to view the results. News
01/03/06 08:00 12M We must have had a power outage over the holiday shutdown period and the 23mmsis receiver compressor failed to restart. As a result the 23mm is warm as of this AM.  

 Copyright Arizona Radio Observatory.
For problems or questions regarding this web contact [tfolkers{at}email{dot}arizona{dot}edu].
Last updated: 02/09/10.